Circular No. 8363

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

UNUSUAL CATACLYSMIC VARIABLE IN HERCULES On June 16, S. Nakano (Sumoto, Japan) reported the discovery by

Yuji Nakamura (Kameyama, Mie-ken, Japan) of a possible nova (mag 11.5) on Tri-X photographs taken on June 13.632 UT with a 200-mm f/4 telephoto lens; the new object was reported by Nakamura to be located at α = 18^{h}39^{m}26^{s}.4, δ = +26^{o}04′09′′ (equinox 2000.0; uncertainties ± 30′′), and he added that nothing was visible around this position on ≈ 50 patrol films taken during 1999–2004 (limiting mag 12). K. Itagaki (Yamagata, Japan, 0.60-m f/5.7 reflector) reports the following position end figures from a CCD image obtained on June 20.488 (with the object at mag 12.8): 26^{s}.16, 09′′.8. J. Bedient (Honolulu, HI) reports that images taken on June 22 with the University of Hawaii 2.2-m reflector suggest that the star in outburst is the western star of a close pair; his position end figures for the variable are 26^{s}.16, 10′′.0, which he notes is 1′′.3 north of 2MASS 18392619+2604087. Following a request by the Central Bureau, T. Matheson, P. Challis, and R. Kirshner (Harvard-Smithsonian Center for Astrophysics) report that a spectrogram (range 370–750 nm), obtained by E. Adams on June 23.32 with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a cataclysmic variable. The spectrum consists of a blue continuum with double-peaked emission lines of hydrogen; the full-width-at-half-maximum of the Hα line is 2000 km/s. T. Harrison and H. Osborne write that a red spectrum (range 700–900 nm) of the object, obtained with the 2.1-m reflector at Kitt Peak on June 22, shows no emission or absorption lines (and so is not a latetype star). A. Price and E. O. Waagen, AAVSO, note that the available photometry of the variable indicate that this is a WZ-Sge-type dwarf nova. Additional magnitudes (CCD and unfiltered unless otherwise noted) for the variable star, provided in part by Waagen and Price: June 17.523, 12.0 (Nakamura; 300-mm f/4 lens + Fuji digital camera); 17.562, 12.3 (Itagaki); 18.138, V = 12.1 (R. James, Las Cruces, NM); 18.309, V = 11.8 (R. Royer, Springville, CA); 19.228, V = 12.3 (James); 20.315, 12.5 (M. Simonsen, Imlay City, MI; visual); 20.488, 12.8 (Itagaki); 21.189, 12.6 (Simonsen); 21.229, V = 12.6 (James); 21.960, 13.0 (G. Poyner, Birmingham, England; visual); 22.050, 13.1 (Poyner); 22.400, V = 12.8 (James); 22.48, 13.2 (Itagaki); 24.044, 13.2 (P. Schmeer, Bischmisheim, Germany; visual); 30.115, 14.3 (M. Mattei, Littleton, MA); 30.156, R = 14.1 (M. Koppelman, Golden Valley, MN); 30.166, V = 14.6 (B. Gary, Hereford, AZ).

2004 June 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green